Michael is 33 times as old as Nadia and is also 16 years older than Nadia. How old is Michael?
Answer by fractalier(6550)

Call their ages, M and N.
From the given facts,
M = 33N and
M = N + 16
Substitute for M and get
33N = N + 16
32N = 16
N = 1/2
M = 16 1/2